OpenStudy (anonymous):

Identify the asymptotes. y=1/x-6

OpenStudy (anonymous):

Horizontal asymptote \[\LARGE y=\frac{1}{x-6} \] \[\LARGE \lim_{x\to\infty }=\frac{1}{x-6} =\frac{\frac 1x}{\frac xx-\frac 6x }=\frac 01=0\] Well dude, I'm not sure, I'll go grab my notebook :$

OpenStudy (anonymous):

Ok

OpenStudy (sburchette):

The function is also undefined at x=6. So there is a vertical asymptote at x=6.

OpenStudy (experimentx):

x = 6 vertical asymptote y = 0 horizontal asymptote

OpenStudy (anonymous):

Horizontal asymptote that I wrote above it's ok. Vertical Asymtote: \[\LARGE y={1\over x-6}\] x-6=0 x=6 and asymptote which is between vertical and horizontal (I STILL don't know how to say it) ... is: \[\LARGE y=kx+b \quad \quad k\neq 0\] \[\LARGE k=\lim_{x\to \infty }\frac{f(x)}{x}\] and \[\LARGE b=\lim_{x\to \infty}[f(x)-kx]\] \[\LARGE k=\lim_{x\to \infty }\frac{1\over x-6 }{x}={1\over x(x-6) }=0\] so since \[\LARGE k\neq 0\] and we got: k=0 there's no asymptote like this :)

OpenStudy (sburchette):

I think you're referring to the oblique asymptote.

OpenStudy (anonymous):

oblique? .. let me go and translate it :F

OpenStudy (anonymous):

whoaa... That's right. At Last. IT'S OBLIQUEEE.... :F